OpenStudy (anonymous):

Use sigma notation to represent the following series for 12 terms. 2- 2/8 + 2/64 - 2/512.....

OpenStudy (anonymous):

\[2-\frac{2}{8}+\frac{2}{64}-\frac{2}{512}+\cdots\\ \frac{2}{8^0}-\frac{2}{8^1}+\frac{2}{8^2}-\frac{2}{8^3}+\cdots\] That should put it in perspective.

OpenStudy (anonymous):

\[(-1)^0\frac{2}{8^0}+(-1)^1\frac{2}{8^0}+(-1)^2\frac{2}{8^0}+(-1)^3\frac{2}{8^0}+\cdots\] Even better